# Learning Content-enhanced Mask Transformer for Domain Generalized Urban-Scene Segmentation

### 09 Dec 2023 v4

@@ Conclusion
In this paper, we explored the feasibility of adapting the mask Transformer for domain-generalized urban-scene semantic segmentation (USSS). To address the challenges of style variation and robust content representation, we proposed a content-enhanced mask attention (CMA) mechanism. This mechanism is designed to capture more resilient content features while being less sensitive to style variations. Furthermore, we extended it to incorporate multi-resolution features and integrate it into a novel framework called the Content-enhanced Mask TransFormer (CMFormer). To evaluate the effectiveness of CMFormer, we conducted extensive experiments on multiple settings. The results demonstrated the superior performance of CMFormer compared to existing domain-generalized USSS methods.
Boarder Social Impact. Given the criticality of safety-related road applications, the proposed method holds significant importance. It has the potential to enhance the accuracy and reliability of semantic segmentation models, thereby contributing to safer and more efficient autonomous systems. Overall, the proposed content-enhanced mask attention mechanism not only offers promising advancements in domain-generalized USSS but also holds potential for broader applications in real-world scenarios.

### 17 Dec 2023 v5

@@ Conclusion
In this paper, we explored the feasibility of adapting the mask Transformer for domain-generalized urban-scene semantic segmentation (USSS). To address the challenges of style variation and robust content representation, we proposed a content-enhanced mask attention (CMA) mechanism. This mechanism is designed to capture more resilient content features while being less sensitive to style variations. Furthermore, we extended it to incorporate multi-resolution features and integrate it into a novel framework called the Content-enhanced Mask TransFormer (CMFormer). Extensive experiments on multiple settings demonstrated the superior performance of CMFormer compared to existing domain-generalized USSS methods.